In ΔRST, angle R = 140° and side s = (3/4)r. Find the measures of angles S and T.
Found 2 solutions by MowMow, ikleyn:
Answer by MowMow(42)

Using the Law of Sines
r/SinR = s/SinS
r/Sin140 = 2/SinS = (3/4)r/SinS
the r's cancel,
so 1/Sin140 =(3/4)/SinS
SinS = (3/4)SinR = (3/4)Sin140 = 0.4821
S = arcSin(0.4821)
S = 11.2 degrees
T = 180 - 140 - 28.8
T = 11.2 degrees

Answer by ikleyn(52788)

.
In ΔRST, angle R = 140° and side s = (3/4)r. Find the measures of angles S and T.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


    The solution in the post by @MowMow is unclear and needs to be polished.
    Below is the solution in the polished form as it should be presented.


Using the Law of Sines

r/sin(140°) = s/sin(S) = ((3/4)r)/sin(S)
the r's cancel,
so 1/sin(140°) =(3/4)/sin(S)
sin(S) = (3/4)*sin(140°) = 0.4821
S = arcsin(0.4821)
S = 28.82°
T = 180° - 140° - 28.82°
T = 11.18°